# Vendi-RAG: Adaptively Trading-Off Diversity And Quality Significantly Improves Retrieval Augmented Generation With LLMs

### v1

@@ 4 Experiments
This section presents a comprehensive evaluation of Vendi-RAG on multi-hop QA datasets. First, we investigate the effectiveness of the Vendi retrieval process in enhancing retrieval diversity. Next, we evaluate the Vendi-RAG pipeline, demonstrating its ability to handle complex queries requiring multi-step reasoning compared to the baselines.
@@
Table 1: Retrieval diversity as measured by the Vendi Score (VS) and Max Pairwise Distance (MPD) for Vendi-RAG, Adaptive-RAG, and Adaptive Retrieval across different datasets. Vendi-RAG achieves higher diversity than the baselines across all datasets and both metrics.
Dataset Adaptive Retrieval Adaptive-RAG Vendi-RAG
VS MPD VS MPD VS MPD
MuSiQue 6.13 1.25 6.55 1.42 7.12 1.95
HotpotQA 4.95 1.10 5.21 1.31 6.82 1.89
2WikiMultiHopQA 5.34 1.32 5.81 1.45 6.68 1.78
Datasets.  Our experiments are conducted on three challenging benchmark multi-hop QA datasets: MuSiQue (Trivedi et al., 2022), HotpotQA (Yang et al., 2018), and 2WikiMultiHopQA (Ho et al., 2020).
@@
Vendi retrieval improves document retrieval diversity.  To assess the impact of the Vendi retrieval process on retrieval diversity, we compared the diversity of outputs from Vendi-RAG against Adaptive-RAG and Adaptive Retrieval. We measured diversity using two different metrics, the VS and the max pairwise distance (MPD). Table 1 summarizes the results. Vendi-RAG achieves higher diversity compared to Adaptive Retrieval and Adaptive-RAG on all dataset, demonstrating its ability to retrieve documents that capture multiple perspectives relevant to the query. This is a crucial advancement, as increased diversity in retrieval directly correlates with improved robustness in multi-hop reasoning (see Table 4). Adaptive-RAG, which incorporates iterative refinement but lacks explicit diversity control, shows moderate retrieval diversity improvement over Adaptive Retrieval.
Accuracy on multi-hop QA tasks.  We further evaluated the performance of the Vendi-RAG pipeline, to assess its ability to reason across multiple documents. The results in Table 4 indicate that Vendi-RAG consistently outperforms other methods in response accuracy across all datasets, showcasing the efficacy of balancing retrieval diversity with quality. Additionally, Vendi-RAG achieves competitive performance in Exact Match (EM) and F1 score. These findings highlight Vendi-RAG’s capability to enhance answer correctness for complex reasoning tasks through improved document retrieval.
Table 2: Performance on multi-hop QA datasets using GPT-3.5 (Turbo). Here we use three different flavors of accuracy: exact match (EM), F1-score (F1), and traditional accuracy (Acc). Vendi-RAG outperforms the baselines in all 3 datasets, except in terms of F1-score, where it performs similarly to Adaptive-RAG.
MuSiQue HotpotQA 2WikiMultiHopQA
Steps Types Methods EM F1 Acc EM F1 Acc EM F1 Acc
Single-step Simple No Retrieval 20.40 31.30 24.40 37.40 51.04 43.20 37.00 48.50 43.40
Single-step Approach 16.40 26.70 23.60 39.60 50.44 45.60 46.80 57.69 52.60
@@
Adaptive-RAG 21.80 32.60 29.60 40.40 52.56 47.00 46.60 60.09 56.80
Vendi-RAG 24.4 32.52 30.4 42.2 57.04 58.4 47.2 58.9 61.4
Impact of the number of retrieved documents on performance.  To further examine the impact of document size on retrieval effectiveness, we analyze the performance of Vendi-RAG and Adaptive-RAG across varying document sizes. Figure 2 illustrates the relationship between document size and performance on the HotPotQA dataset. Vendi-RAG consistently outperforms Adaptive-RAG in accuracy for document sizes greater than two. As document size increases, accuracy improves for both methods, but the gain is notably higher for Vendi-RAG. Similar to accuracy, EM and F1 scores exhibit an increasing trend as document size grows. Vendi-RAG shows a more pronounced improvement, underscoring its capacity to retrieve more informative and relevant documents, thereby enhancing answer quality.
The VS also increases with document size. This is evidence that Vendi-RAG alleviates redundancy since the VS is known to be invariant under duplication (Friedman and Dieng, 2023). An increasing VS indicates less redundancy in the retrieved documents. By leveraging the VS in its retrieval process, Vendi-RAG avoids the redundancy issue that often plagues RAG pipelines.
These results indicate that increasing document size enhances both retrieval diversity and answer correctness. However, the degree of improvement varies across methods, with Vendi-RAG achieving superior gains in all metrics. However, we are computationally bottlenecked primarily by the LLM’s context window limitation and processing time. As the number of retrieved documents increases, we must either truncate documents to fit within the model’s maximum context length or process documents in multiple batches, both of which have significant computational overhead. The bottleneck occurs specifically in the final stage of the pipeline where the LLM processes the retrieved documents to generate answers.
Figure 2: Performance comparison of Vendi-RAG and Adaptive-RAG across different document sizes in terms of Exact Match, F1-score, Accuracy, and Vendi Score on HotPotQA. The backbone LLM here is GPT-4o-mini. Vendi-RAG consistently outperforms Adaptvie-RAG on all the metrics. In particular, performance improves as the number of documents increases.
Performance for different LLM Backbones and retrieval strategies.  To evaluate the impact of different LLM backbones and retrieval strategies on the performance of the Vendi-RAG framework, we conducted experiments using three LLMs: GPT-4o, GPT-4o-mini, and GPT-3.5, across all the multi-hop QA datasets described above. The results, shown in Figure 3, highlight the effectiveness of Vendi-RAG compared to Adaptive-RAG, the best baseline, across all datasets and backbones, except for F1-score on the 2WikiMultiHopQA dataset. In general, larger LLM backbones, such as GPT-4o, achieve superior performance across all datasets, particularly for tasks requiring complex reasoning and synthesis across multiple documents. However, even with smaller models like GPT-4o-mini, the Vendi-RAG model maintains competitive performance, demonstrating its effectiveness.
5 Conclusion
While retrieval-augmented generation (RAG) has proven effective in enhancing large language model (LLM) performance for domain-specific question-answering (QA) tasks, traditional RAG frameworks often struggle with redundancy, particularly in multi-hop reasoning tasks. To address this shortcoming, we introduce Vendi-RAG, a novel framework that jointly optimizes retrieval diversity and answer quality through an iterative refinement process. Vendi-RAG leverages the Vendi Score and an LLM judge to promote semantic diversity while maintaining relevance during retrieval. Our experiments on HotpotQA, MuSiQue, and 2WikiMultiHopQA demonstrate Vendi-RAG’s effectiveness. Specifically, Vendi-RAG outperforms the best baseline by $+4.2\%$ on HotpotQA, $+4.1\%$ on 2WikiMultiHopQA, and $+1.3\%$ on MuSiQue. These gains become even more pronounced as the number of retrieved documents increases, highlighting the importance of retrieval diversity in complex reasoning tasks. Furthermore, we evaluated Vendi-RAG across multiple LLM backbones, including GPT-3.5, GPT-4, and GPT-4o-mini, and observed consistent performance improvements, demonstrating that the framework is model-agnostic. These findings establish Vendi-RAG as an effective and adaptable solution for multi-hop QA.
Limitations.  Vendi-RAG introduces computational overhead due to LLM-based quality scoring, which may limit scalability in real-time applications. Additionally, like all RAG approaches, its performance depends on the quality and completeness of external knowledge sources, making it susceptible to biases or gaps in the retrieved information.
Acknowledgements

### v2

@@ 4 Experiments
In this section, we present a comprehensive evaluation of Vendi-RAG on multi-hop QA tasks. We begin by analyzing the effectiveness of the Vendi retrieval strategy in enhancing retrieval diversity. We then evaluate the full Vendi-RAG pipeline, highlighting its ability to handle complex queries requiring multi-step reasoning, and compare its performance against several strong baselines. All experiments are conducted on three challenging multi-hop QA benchmark datasets: MuSiQue (Trivedi et al., 2022), HotpotQA (Yang et al., 2018), and 2WikiMultiHopQA (Ho et al., 2020) (see Appendix A for additional dataset details).
Dataset Method VS MPD
MuSiQue Adaptive Retrieval 6.13 1.25
Adaptive-RAG 6.55 1.42
Vendi-RAG 7.12 1.95
HotpotQA Adaptive Retrieval 4.95 1.10
Adaptive-RAG 5.21 1.31
Vendi-RAG 6.82 1.89
2WikiMHQA Adaptive Retrieval 5.34 1.32
Adaptive-RAG 5.81 1.45
Vendi-RAG 6.68 1.78
@@
Table 1: Retrieval diversity (Vendi Score (VS) and Max Pairwise Distance (MPD)) across datasets and methods. Vendi-RAG achieves higher diversity than baselines.
Sensitivity Analysis of the VSR Process.
To evaluate the robustness of the VSR process and understand its impact on retrieval diversity, we conducted a sensitivity analysis focusing on how varying the parameter $s$ affects document ranking order within a vector database. This analysis helps elucidate the trade-off between retrieval precision and diversity, which is crucial for enhancing multi-hop reasoning performance. The sensitivity analysis was performed using 100 randomly sampled queries from the dataset to ensure a comprehensive evaluation covering a diverse range of query types and complexity levels. Our primary objective was to investigate how different values of $s$ influence retrieval diversity and ranking consistency.
We evaluated the retrieval pipeline across multiple $s$ values ranging from 0.0 to 1.0, incremented in small steps to capture granular variations in retrieval performance. Setting $s=0.0$ serves as a baseline representing a pure similarity search scenario, where retrieval relies exclusively on cosine similarity or dot product between embeddings, without any emphasis on diversity. This baseline provides a reference point for measuring the impact of increasing $s$ on retrieval diversity. To quantify deviations from the baseline, we employed two complementary ranking comparison metrics:
  • Kendall’s $\tau$: Measures the rank order similarity between two lists by evaluating concordant and discordant pairs. Higher $\tau$ values indicate stronger similarity to the baseline, while lower values reflect greater diversity introduced by increasing $s$.
  • Spearman’s Rank Correlation $\rho$: Assesses the monotonic relationship between two rankings by considering both orderings and positional shifts. Lower $\rho$ values signal substantial deviation from the baseline, indicating increased diversity through higher $s$ values.
The results of the sensitivity analysis are presented in Table 2. As $s$ increases from 0.0 to 1.0, both Kendall’s $\tau$ and Spearman’s $\rho$ decrease progressively, demonstrating that higher $s$ values promote retrieval diversity by prioritizing documents that may be less similar in their embeddings but more relevant from a broader perspective.
Parameter ($s$) $\tau$ $\rho$
0.0 1.00 1.00
0.2 0.797 0.828
0.4 0.688 0.742
0.6 0.485 0.528
0.8 0.265 0.316
1.0 0.074 0.078
Table 2: Sensitivity Analysis of the VSR. Higher $s$ values indicate a greater degree of diversity introduced in the ranking by the retrieval process.
Vendi retrieval improves document retrieval diversity.
To assess the impact of the Vendi retrieval process on retrieval diversity, we compared the diversity of outputs from Vendi-RAG against Adaptive-RAG and Adaptive Retrieval. We measured diversity using two different metrics, the VS and the max pairwise distance (MPD). Table 1 summarizes the results. Vendi-RAG achieves higher diversity compared to Adaptive Retrieval and Adaptive-RAG on all dataset, demonstrating its ability to retrieve documents that capture multiple perspectives relevant to the query. This is a crucial advancement, as increased diversity in retrieval directly correlates with improved robustness in multi-hop reasoning (see Table 3). Adaptive-RAG, which incorporates iterative refinement but lacks explicit diversity control, shows moderate retrieval diversity improvement over Adaptive Retrieval.
MuSiQue HotpotQA 2WikiMultiHopQA
Methods EM F1 Acc EM F1 Acc EM F1 Acc
No Retrieval 20.4 31.3 24.4 37.4 51.0 43.2 37.0 48.5 43.4
Single-step Approach 16.4 26.7 23.6 39.6 50.4 45.6 46.8 57.7 52.6
Adaptive Retrieval 18.8 30.3 24.8 38.6 50.7 43.2 44.2 55.1 50.6
LightRAG 17.4 27.5 23.8 37.4 47.3 42.3 42.2 52.2 47.5
Adaptive-RAG 21.8 32.6 29.6 40.4 52.5 47.0 46.6 60.1 56.8
Self-RAG 1.2 8.2 11.8 5.6 17.8 30.6 3.0 19.1 39.0
Adaptive Retrieval with MMR 22.6 31.0 28.8 41.0 55.0 56.0 45.8 57.0 59.0
Graph-RAG 22.9 32.5 30.1 42.4 57.2 57.8 47.5 59.1 60.9
FiD-Reranker 23.2 33.1 29.8 41.8 56.7 56.9 46.6 59.3 60.1
Vendi-RAG(fixed-$s_{1:N}=0.0$) 22.4 30.4 28.2 40.6 54.1 55.4 45.4 56.8 58.8
Vendi-RAG(fixed-$s_{1:N}=0.3$) 22.4 31.1 28.6 40.7 55.0 55.7 45.3 56.8 59.1
Vendi-RAG(fixed-$s_{1:N}=0.8$) 22.6 30.4 29.1 41.2 55.7 57.2 46.4 57.4 60.2
Vendi-RAG(fixed-$s_{1:N}=0.8$) 23.0 31.2 30.2 42.0 56.9 58.0 47.0 58.7 61.0
Vendi-RAG(fixed-$s_{1:N}=1.0$) 22.9 31.0 29.4 41.6 55.8 57.0 46.4 57.8 60.2
Vendi-RAG($s_{1}=0.8$) 24.4 32.5 30.4 42.2 57.0 58.4 47.2 58.9 61.4
Table 3: Performance on multi-hop QA datasets using GPT-3.5 Turbo is evaluated across three metrics: exact match (EM), F1 score (F1), and traditional accuracy (Acc). Vendi-RAG with $s_{1}=0.8$ outperforms all baselines across the three datasets in terms of EM and Acc, while achieving comparable F1 scores to Adaptive-RAG. Here, Vendi-RAG refers to the variant of Vendi-RAG that excludes the LLM-Judge component.
Accuracy on multi-hop QA tasks.
We further evaluated the performance of the Vendi-RAG pipeline to assess its ability to reason across multiple documents. The results in Table 3 indicate that Vendi-RAG consistently outperforms other methods in response accuracy across all datasets, showcasing the efficacy of balancing retrieval diversity with quality. Additionally, Vendi-RAG achieves competitive performance in Exact Match (EM) and F1 score. These findings highlight Vendi-RAG’s capability to enhance answer correctness for complex reasoning tasks through improved document retrieval. Additionally, we conducted a comprehensive ablation study comparing various retrieval strategies, including Vendi-RAG and MMR, as well as a dynamic adjustment of $s_{t}$ against fixed $s_{t}$ in the table. The results demonstrate that Vendi-RAG with dynamically adjusted $s_{t}$ (with an initial setting of $s_{1}=0.8$) with the LLM-Judge consistently outperforms all baselines, including MMR, across the datasets in terms of EM, F1, and Acc.
Figure 2: Performance comparison of Vendi-RAG and Adaptive-RAG across different document sizes in terms of Exact Match, F1-score, Accuracy, and Vendi Score on HotPotQA. The backbone LLM used is GPT-4o-mini. Vendi-RAG consistently outperforms Adaptive-RAG across all metrics. In particular, performance improves as the number of retrieved documents increases. Different variants of Vendi-RAG are plotted based on the fixed initialization value $s_{1}$ for the diversity-relevance parameter $s_{t}$, with $s_{1}=0.8$ achieving the best overall results.
Ablation Study on Retrieval Strategy and $s_{t}$ Scheduling.
To better understand the effectiveness of our retrieval strategy, we conducted a comprehensive ablation study, examining various configurations of the Vendi-RAG pipeline. First, we evaluated fixed values of $s_{t}$ across all steps, testing settings such as $s_{t}=\{0.0,0.3,0.8,1.0\}$. Among the fixed schedules, setting $s_{t}=0.8$ consistently achieved the best overall performance in terms of Exact Match (EM), F1 score, and Accuracy across all datasets, highlighting the importance of dynamically balancing retrieval diversity with quality. We further compared Vendi-RAG against traditional retrieval baselines, including Adaptive Retrieval with MMR, and observed that our method outperformed the MMR retrieval method across all metrics. Additionally, we tested a variant without dynamic scheduling (fixed $s_{t}$) and a variant without the LLM-Judge module (Vendi-RAG). The results show that dynamically adjusting $s_{t}$ during retrieval using the LLM-Judge significantly boosts performance compared to fixed schedules and simpler retrieval strategies. These findings emphasize the critical role of adaptive retrieval and document assessment in enabling Vendi-RAG to effectively handle complex multi-hop reasoning tasks.
Impact of the number of retrieved documents on performance.
To further examine the impact of document size on retrieval effectiveness, we analyze the performance of Vendi-RAG and Adaptive-RAG across varying document sizes and initial settings of $s_{1}=\{0.3,0.8,1.0\}$. Figure 2 illustrates the relationship between document size and performance on the HotPotQA dataset. Vendi-RAG consistently outperforms Adaptive-RAG in accuracy for document sizes greater than two with any $s_{1}$ setting. As document size increases, accuracy improves for both methods, but the gain is notably higher for Vendi-RAG. Similar to accuracy, EM and F1 scores exhibit an increasing trend as document size grows. Vendi-RAG shows a more pronounced improvement, underscoring its capacity to retrieve more informative and relevant documents, thereby enhancing answer quality. The VS also increases with document size. This is evidence that Vendi-RAG alleviates redundancy since the VS is known to be invariant under duplication (Friedman and Dieng, 2023). An increasing VS indicates less redundancy in the retrieved documents. By leveraging the VS in its retrieval process, Vendi-RAG avoids the redundancy issue that often plagues RAG pipelines. These results indicate that increasing document size enhances both retrieval diversity and answer correctness. Vendi-RAG is achieving superior gains in all metrics. However, we are computationally bottlenecked primarily by the LLM’s context window limitation and processing time. As the number of retrieved documents increases, we must either truncate documents to fit within the model’s maximum context length or process documents in multiple batches, both of which have significant computational overhead.
Performance for different LLM Backbones and retrieval strategies.
To evaluate the impact of different LLM backbones and retrieval strategies on the performance of the Vendi-RAG (with $s_{1}=0.8$) framework, we conducted experiments using three LLMs: GPT-4o, GPT-4o-mini, and GPT-3.5, across all the multi-hop QA datasets described above. The results, shown in Figure 3, highlight the effectiveness of Vendi-RAG compared to Adaptive-RAG, the best baseline, across all datasets and LLM-backbones, except for F1-score on the 2WikiMultiHopQA dataset. In general, larger LLM backbones, such as GPT-4o, achieve superior performance across all datasets, particularly for tasks requiring complex reasoning and synthesis across multiple documents. However, even with smaller models like GPT-4o-mini, the Vendi-RAG model maintains competitive performance, demonstrating its effectiveness.
VSR vs MMR.
VS offers key advantages over traditional diversity metrics like MMR. While MMR relies on pairwise comparisons to balance relevance and novelty, it lacks a global view of semantic diversity across the retrieved set. In contrast, VS is a principled, global metric based on the eigenvalues of the normalized kernel matrix, directly measuring the effective number of distinct documents. It reaches its maximum when documents are entirely unique and its minimum when they are identical, providing an intuitive, mathematically grounded measure of diversity. This global perspective makes VS particularly effective for multi-hop QA, where broad semantic coverage is critical. Moreover, VS integrates naturally with Vendi-RAG’s dynamic retrieval adjustment, enabling fine-grained control over the diversity-relevance trade-off via a single parameter, and addressing the challenge of balancing coverage and precision in complex reasoning tasks.
5 Conclusion
While RAG has proven effective in enhancing LLM performance for domain-specific QA tasks, traditional models often struggle with redundancy, particularly in multi-hop reasoning tasks. To address this shortcoming, we introduce Vendi-RAG, a novel framework that jointly optimizes retrieval diversity and answer quality through an iterative refinement process. Vendi-RAG leverages the Vendi Score and an LLM judge to promote semantic diversity while maintaining relevance during retrieval. Our experiments on HotpotQA, MuSiQue, and 2WikiMultiHopQA demonstrate Vendi-RAG’s effectiveness. Specifically, Vendi-RAG outperforms the best baseline by $+4.2\%$ on HotpotQA, $+4.1\%$ on 2WikiMultiHopQA, and $+1.3\%$ on MuSiQue. These gains become even more pronounced as the number of retrieved documents increases, highlighting the importance of retrieval diversity in complex reasoning tasks. Furthermore, we evaluated Vendi-RAG across multiple LLM backbones, including GPT-3.5, GPT-4, and GPT-4o-mini, and observed consistent performance improvements, demonstrating that the framework is model-agnostic. These findings establish Vendi-RAG as an effective and adaptable solution for multi-hop QA.
6 Limitations
Vendi-RAG introduces computational overhead due to LLM-based quality scoring, which may limit scalability in real-time applications. Additionally, like all RAG approaches, its performance depends on the quality and completeness of external knowledge sources, making it susceptible to biases or gaps in the retrieved information.
7 Ethics Statement
The deployment of LLMs, including their use in Vendi-RAG, necessitates careful ethical consideration. Since the model relies on external knowledge sources, concerns arise regarding the credibility and accuracy of retrieved content. Ensuring the reliability and factual integrity of information is crucial to mitigating risks related to bias and misinformation.
Acknowledgements
Appendix A Datasets
@@
Appendix B Evaluation Metrics
To compare model performance across different datasets, we employ the following key evaluation metrics:
  • Exact Match (EM): This metric calculates the percentage of predictions that exactly match the ground truth answers. It is defined as:
    EM=Number of exact matchesTotal number of $\text{EM}=\frac{\text{Number of exact matches}}{\text{Total number of queries}}\times 100$ (4)
    EM is a strict metric that grants credit only when the predicted answer matches the ground truth exactly in both content and format. It is particularly useful for assessing a model’s precision in generating accurate responses.
  • F1 Score (F1): The F1 score captures the harmonic mean of precision and recall at the token level, providing a balanced measure of correctness. It is defined as:
@@ Appendix B Evaluation Metrics
    where precision is the fraction of retrieved tokens that are relevant, and recall is the fraction of relevant tokens that are retrieved. The F1 score is particularly relevant for multi-hop QA tasks, where partial correctness (e.g., retrieving some but not all supporting evidence) is informative.
  • Accuracy (Acc): Accuracy measures the proportion of correct predictions over all evaluated queries. It is defined as:
    Acc=Number of correct predictionsTotal number of $\text{Acc}=\frac{\text{Number of correct predictions}}{\text{Total number of queries}}\times 100$ (6)
    Unlike EM, which requires exact matches, accuracy provides a broader assessment by capturing overall correctness, including responses that convey the intended meaning.
  • Max Pairwise Distance (MPD): This metric evaluates the maximum Euclidean distance between pairs of retrieved data points, measuring diversity. It is defined as:
    $\text{MPD}=\max_{i,j}\|x_{i}-x_{j}\|_{2},\quad i\neq j$ (7)
    where $x_{i}$ and $x_{j}$ represent document embeddings in the feature space. Higher values indicate greater diversity among retrieved documents.
Each of these metrics offers a unique perspective on model performance. EM is a stringent measure of precision, F1 balances precision and recall, and accuracy provides an overall correctness measure. Meanwhile, MPD and diversity-based metrics assess the variety and independence of retrieved documents, critical for multi-hop QA tasks requiring integration of diverse information.
Appendix C Implementation Details
The Vendi-RAG framework employs dense vector representations derived from transformer-based embeddings to compute the similarity score $\text{SS}(q,\mathcal{D})$ between a query $q$ and a set of documents $\mathcal{D}$. This high-dimensional semantic comparison allows the model to effectively capture contextual relationships and retrieve relevant documents across diverse domains. To evaluate answer quality, we use a consistent LLM backbone with a specifically designed prompt that positions the LLM as an expert judge. This judge-based evaluation framework assesses the generated answers according to this prompt:
LLM as a Judge Prompt You are an expert LLM-based judge tasked with evaluating the quality of answers in a Retrieval-Augmented Generation (RAG) system. Your evaluation will consider the following aspects:
1. Coherence: Assess whether the provided answer is logically consistent and flows smoothly, without conflicting statements or gaps in reasoning. 2. Relevance: Evaluate how well the answer addresses the query based on the information from the retrieved documents. 3. Query Alignment: Determine how closely the answer aligns with the specific query asked, ensuring that the response is focused and appropriate. Your evaluation will be quantified based on the following scoring system:
- Coherence Score (C): [1 - 10], where 10 is perfectly coherent.
- Relevance Score (R): [1 - 10], where 10 is highly relevant to the query. - Query Alignment Score (Q): [1 - 10], where 10 is perfectly aligned. Provide a quality score $Q_{t}$ as the average of these individual scores: $Q_{t}=\text{mean}(C,R,Q)$ Query: {query}
This judge assesses coherence, relevance, and alignment with the query to produce a quality score $Q_{t}$. The quality threshold ($\tau$) is set to 0.85 for all experiments, ensuring a consistent standard of answer evaluation. While we initially set $s_{1}=0.8$ across experiments to prioritize diversity, our ablation study explores this hyperparameter through testing different fixed values and employing dynamic adjustment. This analysis highlights the importance of dynamic adjustment for improved performance across diverse datasets.
C.1 Dataset Processing and Chunking
Preparing datasets for question-answering requires transforming data into a searchable vector database to enable efficient retrieval. This workflow includes document chunking and semantic embedding to optimize performance.
The dataset, provided in JSON format with context paragraphs and metadata, is processed by splitting each document into smaller chunks. Each chunk has a maximum size of 512 tokens, with a 50-token overlap to preserve context across chunk boundaries and facilitate multi-hop reasoning in long documents.
C.2 Embedding Model and Vector Database
We use the SentenceTransformer model, specifically all-mpnet-base-v2, to generate dense vector representations for documents and queries. These embeddings are stored locally to avoid redundant downloads and improve reusability. The Chroma vector database efficiently stores and retrieves these vectorized documents along with metadata, such as document titles and chunk IDs.
C.3 Batch Processing and Database Population
To efficiently populate the vector database, document chunks are processed in batches of up to 10,000. This approach optimizes memory usage while ensuring completeness in the ingestion process. The total number of processed chunks is logged for verification.
C.4 Query Answering Workflow
For queries such as "Who is the father-in-law of Queen Hyojeong?", relevant chunks are retrieved using Chroma’s similarity-based search mechanism. The system ranks the top 10 chunks based on their semantic similarity to the query, leveraging embeddings generated by all-mpnet-base-v2 to ensure precise and relevant results.
C.5 Key Configuration Details
The system is configured with the following parameters:
  • Embedding Model: all-mpnet-base-v2, optimized for semantic similarity.
  • Vector Database: Chroma, persisted to disk for efficient reuse.
  • Chunk Size: 512 tokens per chunk, with a 50-token overlap.
  • Batch Size: Up to 10,000 chunks per batch.